# LubDubDecoder: Bringing Micro-Mechanical Cardiac Monitoring to Hearables

### v1

@@ Abstract.
We present LubDubDecoder, a system that enables fine-grained monitoring of micro-cardiac vibrations associated with the opening and closing of heart valves across a range of hearables. Our system transforms the built-in speaker, the only transducer common to all hearables, into an acoustic sensor that captures the coarse “lub-dub” heart sounds, leverages their shared temporal and spectral structure to reconstruct the subtle seismocardiography (SCG) and gyrocardiography (GCG) waveforms, and extract the timing of key micro-cardiac events. In an IRB-approved feasibility study with 18 users, our system achieves correlations of 0.88–0.95 compared to chest-mounted reference measurements in within-user and cross-user evaluations, and generalizes to unseen hearables using a zero-effort adaptation scheme with a correlation of 0.91. Our system is robust across remounting sessions and music playback.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spain
Figure 1. LubDubDecoder enables hands-free monitoring of micro-mechanical cardiac events across a range of hearables, including (a) an older adult sleeping at home with over-ear headphones, (b) a driver wearing wireless earbuds, and (c) a commuter on a train using bone-conduction earphones. points correspond to micro-cardiac events detected by our system. (d) Smartphone app user interface with hearables.
[…]
This figure is composed of four sub-figures that collectively illustrate the system’s ability to monitor micro-mechanical cardiac events using different types of hearable devices in various everyday settings. Sub-figures ’a’, ’b’, and ’c’ are illustrations, while sub-figure ’d’ is a photograph. Sub-figure ’a’ shows a person wearing over-ear headphones and sleeping on a couch, with a red speech bubble above their head containing a red waveform that represents the captured cardiac signal. Sub-figure ’b’ shows a person in a car wearing a green wireless earbud while driving, with a green speech bubble above their head containing a green waveform representing a captured signal. Sub-figure ’c’ shows a person on a train reading a book while wearing a purple bone-conduction earphone, with a purple speech bubble above their head containing a purple waveform. In all three illustrations, blue circular markers on the waveforms indicate detected micro-cardiac events. Sub-figure ’d’ is a photograph of the system’s components: a smartphone with the user interface of the mobile app is shown in the center, displaying a user profile, the original hearable recording, and the reconstructed SCG and GCG signals. Surrounding the smartphone are four types of hearables compatible with the system: a pair of black over-ear headphones at the top, a pair of white wireless earbuds to the left, a pair of white wired earphones to the right, and a black bone-conduction headphone to the right.
@@ 1. Introduction
Figure 2. System overview. LubDubDecoder reconstructs the heart’s micro-mechanical signals from coarse-grained heart sounds recorded at the ear using microphones and speakers across a range of hearables. It identifies the timing of key micro-cardiac events, and enables hands-free monitoring of cardiovascular health in everyday scenarios.
[…]
Hearables, by contrast, occupy a fixed anatomical position in the ear, minimizing placement variability across removal and re-mounting sessions. Given that micro-cardiac vibrations propagate from the chest to the ear canal, hearables present themselves as a promising platform for convenient everyday monitoring of these signals (Fig. 1). Imagine tracking fine-grained cardiac activity passively as an older adult sleeps at home using over-ear headphones. Or envision hands-free monitoring on earbuds during everyday routines such as working at a desk or driving where cardiac abnormalities may go unnoticed. Even during a daily commute, bone-conduction earphones could capture these signals unobtrusively.
We present LubDubDecoder, the first hearable system that bring micro-mechanical cardiac monitoring out of the clinic and enables monitoring in everyday scenarios across a range hearables (Fig. 2). While prior work (Fu et al., 2025) has reconstructed SCG using IMU-equipped hearables, IMUs are typically only found on true wireless stereo earbuds (Fu et al., 2025) or higher-end earphones such as those supporting active noise cancellation (Chen et al., 2024). Other approaches measure SCG using IMUs on smartphones (Wang et al., 2018; Lee et al., 2021; Hossein et al., 2024) and contactless mmWave radar (Ha et al., 2020), however, these require strapping or holding the phone against the chests, or constraining users to remain in a fixed location and orientation. In contrast, our approach is broadly compatible across hearables and enables monitoring across environments, including during music playback.
@@
    To achieve this, we transform the speaker, the only transducer common to all hearables, into a microphone to capture and amplify the soft, coarse-grained heart sounds from the ear and reconstruct the SCG and GCG signal. Our system works across hearables with different designs including bone-conduction earphones that contact the outer ear, over-ear headphones that enclose the ears, as well as both wired and wireless earbuds.
@@
Below we summarize the main contributions of our system:
  • LubDubDecoder reconstructs the fine-grained SCG and GCG waveforms from ear-based heart sounds and extracts the timing of key micro-cardiac events in an IRB-approved feasibility study with 18 subjects, achieving Pearson correlations between 0.88–0.95 compared to conventional chest-mounted measurements in both within-user and cross-user evaluations.
@@
  • We perform a user experience survey demonstrating high levels of ease of use (4.5 of 5) and system trustworthiness (4.1 of 5).
@@ 2. Background and related work
Prior work on cardiac monitoring with mobile devices. While prior work on hearables have been designed for cardiac monitoring  (Cao et al., 2024; Butkow et al., 2023; Fan et al., 2023; Li et al., 2024; Chen et al., 2024; Fan et al., 2021; Gilliam III et al., 2022) they are mostly focused on reconstructing PCG, ECG and audioplethysmography (APG, which is similar to PPG) signals from in-ear cardiac sounds, we present the first system to reconstruct micro-mechanical cardiac signals across a range of hearables, from ear-based cardiac sounds and extract the timing of micro-cardiac events.
While prior work (Fu et al., 2025) has focused on leveraging the IMU on hearables to reconstruct SCG, these sensors only sample at 20-25 Hz and suffer from packet loss since audio streaming is prioritized. As a result, super-resolution algorithms are needed to interpolate the signal, but it is challenging to reliably capture the timing of micro-cardiac events, given that SCG contains frequency components up to 50 Hz and thus requires sampling at 100 Hz. Furthermore IMUs are a specialized sensor typically only present on true wireless stereo earbuds (Fu et al., 2025) or higher-end earphones (Chen et al., 2024).
@@
Radar-based systems (Ha et al., 2020) have leveraged mmWave to capture chest vibrations corresponding to SCG. This approach requires users to remain directly in front of the radar within an operational range of 25–50 cm, and relies on comparatively expensive devices ($\approx\$350$ (ti_, 2025)). Because the radar is typically mounted in a fixed position, it requires users to carry and reposition it wherever they want to take a measurement. In contrast, our approach is inherently mobile and can be worn throughout daily life across different environments, and with greater flexibility on user position.
@@ 3.2.2. Measuring waveform variation
Ear-based cardiac sounds SCG GCG
Intra-session 0.87 $\pm$ 0.11 0.86 $\pm$ 0.07 0.87 $\pm$ 0.06
Inter-session 0.82 $\pm$ 0.09 0.73 $\pm$ 0.19 0.79 $\pm$ 0.07
Inter-user 0.31 $\pm$ 0.19 0.03 $\pm$ 0.24 -0.03 $\pm$ 0.36
Inter-device 0.15 $\pm$ 0.36 0.22 $\pm$ 0.19 0.42 $\pm$ 0.29
Table 2. Cardiac waveform variability across different experimental conditions. Waveform variability is measured using Pearson correlation coefficient, with mean and standard deviation reported. Ear-based cardiac sounds are recorded from the in-ear speaker of earphones. The SCG and GCG signal were recorded from the smartphone IMU placed against the chest.
Intra-session Inter-session Inter-user Inter-device
SCG AO–MC (ms) 2.6 6.7 29.3 31.2
AO–IM (ms) 1.4 2.6 18.8 19.7
AO–MA (ms) 5.0 3.5 13.1 14.6
AO–RE (ms) 7.6 9.1 26.2 27.6
Intra-session Inter-session Inter-user Inter-device
GCG AO–MC (ms) 2.2 8.5 27.8 31.7
AO–IM (ms) 3.9 5.5 13.8 18.4
AO–MA (ms) 6.8 1.7 24.8 12.4
AO–RE (ms) 7.3 7.2 29.7 23.9
Table 3. Timing variability of fiducial point timing across different experimental conditions. Standard deviations for the timing between the aortic opening (AO) and four additional fiducial points in the SCG and GCG signal. The SCG and GCG signal were recorded from the smartphone IMU placed against the chest.
Intra-session. This analysis provides a baseline for cardiac cycle variability over time within an individual. To characterize this, we had a single participant wear the wireless earbuds and place the smartphone on their chest without removal or re-mounting over the course of 2 minutes. We then segmented the ear-based cardiac sounds, SCG, and GCG signals into individual cardiac cycles of fixed duration (800 ms) corresponding to the typical length of a single cardiac cycle (Widmaier et al., 2022) (further description of algorithm in Sec. 3.3.2).
@@
[…]
Inter-session. We next consider how much the signals change when the device is removed and re-mounted. This is an important practical consideration as hearables are frequently taken off and worn again in daily life. To evaluate this, we recruited two participants as before and perform seven remounting trials. In each trial, the hearable was removed and re-mounted, as well as the smartphone for SCG and GCG data collection, followed by a 60-second recording session.
@@
[…]
Inter-user. Beyond temporal and remounting variability, another critical factor is how cardiac signals vary across different individuals. Inter-user variability reflects intrinsic physiological differences that affect how hearables and smartphones capture the signals. To assess this, we analyzed recordings collected from 6 randomly selected subjects from our human subjects study using the same hearable device (in-ear mic from the wireless earbuds). Details of the population are described in Sec. 4.2.
@@
To evaluate device-related variability, we recorded ear-based cardiac sounds from three participants, each using five different hearable devices (Fig. 5), and the SCG and GCG signal across four different smartphones (iPhone 12, Motorola Edge, Google Pixel 7, and Samsung Galaxy S9). We show the variation in the physiological signals for one of the participants in Fig. 8 which illustrates that the ear-based cardiac sounds exhibited substantial differences across devices. We then show the quantitative variation for this inter-device variation averaged across three users and show the Pearson correlation for the ear-based cardiac sounds, SCG, and GCG signal was 0.15 $\pm$ 0.36, 0.22 $\pm$ 0.19, and 0.42 $\pm$ 0.29 respectively (Table 2), which is lower than the intra-session and inter-session similarities. The fiducial point timing variability is also higher, ranging from 12 to 31 ms for the SCG and GCG signals, in comparison to the intra-session variability of 1 to 8 ms (Table 3).
@@ 3.3.1. Motion artifact removal pipeline
We recorded a dataset containing six types of motion events: brow raising, chewing, nodding, talking, and walking. All recordings were collected from a single subject using in-ear microphones. For each acoustic event, we recorded 10-second segments, along with a set of recordings when the user was static and there was no external interference or music playback (Fig. 10a). Fig. 10b shows that the various motion types produced distinct signal distortions that were clearly visible and would obscure the smaller heart sounds.

### v2

@@ Abstract.
We present LubDubDecoder, a system that enables fine-grained monitoring of micro-cardiac vibrations associated with the opening and closing of heart valves across a range of hearables. Our system transforms the built-in speaker, the only transducer common to all hearables, into an acoustic sensor that captures the coarse “lub-dub” heart sounds, leverages their shared temporal and spectral structure to reconstruct the subtle seismocardiography (SCG) and gyrocardiography (GCG) waveforms, and extract the timing of key micro-cardiac events. In an IRB-approved feasibility study with 25 users, our system achieves correlations of 0.88–0.95 compared to chest-mounted reference measurements in within-user and cross-user evaluations, and generalizes to unseen hearables using a zero-effort adaptation scheme with a correlation of 0.91. Our system is robust across remounting sessions and music playback.
Earables, Cardiovascular health, Wearable health monitoring, Well-being, Seismocardiography
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790445isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Ubiquitous and mobile computing systems and tools
Figure 1. LubDubDecoderSystem overview. (a) LubDubDecoder reconstructs the heart’s micro-mechanical signals from coarse-grained heart sounds recorded at the ear using microphones and speakers across a range of hearables. It identifies the timing of key micro-cardiac events, and enables hands-free monitoring of cardiovascular health in everyday scenarios. (b) Smartphone app user interface with hearables.
@@ 1. Introduction
Hearables, by contrast, occupy a fixed anatomical position in the ear, minimizing placement variability across removal and re-mounting sessions. Given that micro-cardiac vibrations propagate from the chest to the ear canal, hearables present themselves as a promising platform for convenient everyday monitoring of these signals.
We present LubDubDecoder, the first hearable system that bring micro-mechanical cardiac monitoring out of the clinic and enables monitoring in everyday scenarios across a range hearables (Fig. 1). While prior work (Fu et al., 2025) has reconstructed SCG using IMU-equipped hearables, IMUs are typically only found on true wireless stereo earbuds (Fu et al., 2025) or higher-end earphones such as those supporting active noise cancellation (Chen et al., 2024). Other approaches measure SCG using IMUs on smartphones (Wang et al., 2018; Lee et al., 2021; Hossein et al., 2024) and contactless mmWave radar (Ha et al., 2020), however, these require strapping or holding the phone against the chests, or constraining users to remain in a fixed location and orientation. In contrast, our approach is broadly compatible across hearables including low-cost earphones and enables monitoring across environments, including during music playback.
@@
    To achieve this, we transform the speaker, the only transducer common to all hearables, into a microphone to capture and amplify the soft, coarse-grained heart sounds from the ear and reconstruct the SCG and GCG signal. Our system works across hearables with different designs and varied price points including bone-conduction earphones that contact the outer ear, over-ear headphones that enclose the ears, as well as both wired and wireless earbuds.
@@
This work contributes to the HCI community through:
  • A novel sensing approach that repurposes the in-ear speakers for cardiac sensing, transforming everyday hearables into an equitable health monitoring platform that remains accessible to individuals with limited income or low technology literacy.
  • An automated labeling pipeline for high-throughput segmentation and annotation of cardiac signals that can be applied to other cardiac health monitoring contexts.
  • Experimental protocols and a motion artifact detection algorithm for assessing feasibility in free living deployments across everyday static and dynamic activities and during music playback, which other researchers can build on.
Below we summarize the main technical contributions of our system:
  • LubDubDecoder reconstructs the fine-grained SCG and GCG waveforms from ear-based heart sounds and extracts the timing of key micro-cardiac events in an IRB-approved feasibility study with 25 subjects, achieving Pearson correlations between 0.88–0.95 compared to conventional chest-mounted measurements in both within-user and cross-user evaluations.
@@
  • We perform a user experience survey demonstrating high levels of ease of use (4.6 of 5) and system trustworthiness (4.3 of 5).
Contribution to hearables and cardiac sensing research in HCI. Conceptually, to the best of our knowledge, we are the first HCI system that is focused around using the built-in speaker, the only transducer common to all hearables, to perform cardiac monitoring across broader range of hearables of varying price points. In contrast, prior works use dedicated IMUs or in-ear microphones which are not present on most classes of hearables in particular low-cost earphones (Fu et al., 2025). We show that despite weaker acoustic output, speaker-based sensing can enable robust cardiac monitoring at comparable performance to in-ear mics (Fig. 13). This specific design choice is what allows us to have the benefits of greater ubiquity without trading off system performance.
We focus on extracting SCG/GCG signals as it is unique amongst cardiac signals in its ability to measure sub-audible mechanical events not measurable by PCG and can assess cardiac conditions (Iftikhar et al., 2018; Pandia et al., 2012; Pänkäälä et al., 2016; Sahoo et al., 2018; Salerno and Zanetti, 1991; Solar et al., 2017; Tadi et al., 2016; Taebi and Mansy, 2017; Tavakolian, 2010; Tavakolian et al., 2010; Zakeri et al., 2016), and in some cases even outperforming ECG electrical signals (Wilson et al., 1993). This particular set of cardiac signals is under-explored in comparison to prior works which have focused on extracting ECG (Butkow et al., 2023), PCG (Chen et al., 2024), respiration volume monitoring (Liu et al., 2025), and arterial pressure waveform (Christofferson et al., 2025).
Enabling practical SCG/GCG monitoring requires contributions from a system design perspective:
  1. (1)
    Cross condition characterization. We perform a detailed characterization of micro-cardiac waveform variability in Sec. 3 for intra-session, inter-session, inter-user and inter-device scenarios which motivates the design of our system, and reveals (a) the susceptibility of conventional smartphone-based SCG/GCG monitoring to placement variability (b) the comparative stability of in-ear SCG/GCG signals even after remounting (c) the need for per-person and per-device calibration.
  2. (2)
    Novel calibration framework. Based on these findings, we introduce in Sec. 3.4.5 (i) a human-in-the-loop adaptation procedure for new individuals which incorporates a brief simultaneous measurement using the smartphone on the chest and the hearable, and (ii) a zero-effort normalization strategy for new hearables without any explicit effort on the part of the user that normalizes the frequency response of a new device with recordings from a previously seen device.
  3. (3)
    Micro-cardiac event timing reconstruction. With this distinct setup design and formulation, we are able to reconstruct the timing of fiducial points from the ear in cross-session, cross-user and cross-device scenarios (Sec. 5.3) which is not shown in prior works (Fu et al., 2025).
Ethics and consent. This work includes (1) a feasibility study evaluating the cardiac waveform reconstruction performance of our system, (2) a post study user experience survey, and (3) recordings to characterize the cardiac signals and train a motion artifact detection classifier. For each of these components, informed written consent was obtained prior to participation and prior to any data collection. All these procedures were conducted under IRB approval (STUDY2025_00000155) and complied with relevant ethical regulations.
@@ 2. Background and related work
Islam et al. (2025) (Islam et al., 2025) Hearable IMU and PPG sensor
Liu et al. (2025) (Liu et al., 2025) Hearable mic
Christofferson et al. (2025) (Christofferson et al., 2025) Hearable mic
@@
Prior work on cardiac monitoring with mobile devices. While prior work on hearables have been designed for cardiac monitoring  (Cao et al., 2024; Butkow et al., 2023; Fan et al., 2023; Li et al., 2024; Chen et al., 2024; Fan et al., 2021; Gilliam III et al., 2022; Liu et al., 2025; Christofferson et al., 2025) they are mostly focused on reconstructing PCG, ECG and audioplethysmography (APG, which is similar to PPG) signals, arterial pressure waveforms, and respiration volume from in-ear cardiac sounds, we present the first system to reconstruct micro-mechanical cardiac signals across a range of hearables, from ear-based cardiac sounds and extract the timing of micro-cardiac events.
Recent IMU-based approaches including BallistoBud (Islam et al., 2025) and TWSCardio (Fu et al., 2025) have demonstrated the feasibility of capturing cardiac motion using earbud accelerometers. A limitation of this approach is that IMUs are specialized sensors found mainly in true wireless stereo earbuds (Fu et al., 2025) or high-end hearables (Chen et al., 2024). In contrast, our approach reconstructs fine-grained cardiac waveforms from widely available in-ear microphones and speakers, which is an approach that applies more broadly across even low-end hearables.
@@
Radar-based systems (Ha et al., 2020) have leveraged mmWave to capture chest vibrations corresponding to SCG. This approach requires users to remain directly in front of the radar within an operational range of 25–50 cm, and relies on comparatively expensive devices ($\approx\mathdollar 350$ (ti_, 2025)). Because the radar is typically mounted in a fixed position, it requires users to carry and reposition it wherever they want to take a measurement. In contrast, our approach is inherently mobile and can be worn throughout daily life across different environments, and with greater flexibility on user position.
@@ 3. System design
3.1. Designing for accessible, everyday use
Our design was guided by principles of accessibility, comort, and everyday usability. We prioritized a hardware-agnostic approach that works across a wide range of hearables including low-end ones to reduce adoption barriers and cost. Leveraging the built-in speaker and microphone enables passive monitoring without requiring users to wear dedicated sensors or require active user participation, which lower the barrier to adoption particularly for older adults who may have cognitive impairment or have lowered technology literacy rates. The ear location was chosen for its consistent coupling and comfort during long-term wear. These considerations align with prior HCI work emphasizing unobtrusive, low-effort health monitoring and sustainable engagement (Karkar et al., 2017; Mariakakis et al., 2018).
@@ 3.3.1. Hearables for capturing cardiac sounds
Ear sounds SCG GCG
Intra-session 0.88 $\pm$ 0.08 0.89 $\pm$ 0.06 0.87 $\pm$ 0.06
Inter-session 0.82 $\pm$ 0.09 0.73 $\pm$ 0.19 0.79 $\pm$ 0.07
Inter-user 0.31 $\pm$ 0.19 0.03 $\pm$ 0.24 -0.03 $\pm$ 0.36
Inter-device 0.15 $\pm$ 0.36 0.22 $\pm$ 0.19 0.42 $\pm$ 0.29
Table 2. Cardiac waveform variability across different experimental conditions. Waveform variability is measured using Pearson correlation coefficient (mean $\pm$ std). Ear sounds are recorded from the in-ear speaker. SCG and GCG are recorded from the smartphone IMU against the chest.
Intra
session
Inter
session
Inter
user
Inter
device
SCG AO–MC 2.6 6.7 29.3 31.2
AO–IM 1.4 2.6 18.8 19.7
AO–MA 5.0 3.5 13.1 14.6
AO–RE 7.6 9.1 26.2 27.6
GCG AO–MC 2.2 8.5 27.8 31.7
AO–IM 3.9 5.5 13.8 18.4
AO–MA 6.8 1.7 24.8 12.4
AO–RE 7.3 7.2 29.7 23.9
Table 3. Timing variability of fiducial point timing across different experimental conditions. Standard deviations (ms) for timing from aortic opening (AO) to four fiducial points in SCG and GCG from the smartphone IMU against the chest.
@@ 3.3.2. Measuring waveform variation
Intra-session. This analysis provides a baseline for cardiac cycle variability over time within an individual. To characterize this, we had a single participant wear the wireless earbuds and place the smartphone on their chest without removal or re-mounting over the course of 2 minutes. We then segmented the ear-based cardiac sounds, SCG, and GCG signals into individual cardiac cycles of fixed duration (800 ms) corresponding to the typical length of a single cardiac cycle (Widmaier et al., 2022) (further description of algorithm in Sec. 3.4.2). We note that this participant was a member of the study team that provided informed written consent prior to any data collection.
@@
Inter-session. We next consider how much the signals change when the device is removed and re-mounted. This is an important practical consideration as hearables are frequently taken off and worn again in daily life. To evaluate this, we recruited two participants as before and perform seven remounting trials. In each trial, the hearable was removed and re-mounted, as well as the smartphone for SCG and GCG data collection, followed by a 60-second recording session. We note that these two participants are members of the study team that provided informed written consent prior to any data collection.
@@
Inter-user. Beyond temporal and remounting variability, another critical factor is how cardiac signals vary across different individuals. Inter-user variability reflects intrinsic physiological differences that affect how hearables and smartphones capture the signals. To assess this, we analyzed recordings collected from 6 randomly selected subjects from our human subjects study using the same hearable device (in-ear mic from the wireless earbuds). We describe in Sec. 4.2 details of the population and note that all participants provided informed written consent prior to any data collection.
@@
To evaluate device-related variability, we recorded ear-based cardiac sounds from three participants, each using five different hearable devices (Fig. 4), and the SCG and GCG signal across four different smartphones (iPhone 12, Motorola Edge, Google Pixel 7, and Samsung Galaxy S9). We note that these three participants are members of the study team that provided informed written consent prior to any data collection. We show the variation in the physiological signals for one of the participants in Fig. 7 which illustrates that the ear-based cardiac sounds exhibited substantial differences across devices. We then show the quantitative variation for this inter-device variation averaged across three users and show the Pearson correlation for the ear-based cardiac sounds, SCG, and GCG signal was 0.15 $\pm$ 0.36, 0.22 $\pm$ 0.19, and 0.42 $\pm$ 0.29 respectively (Table 2), which is lower than the intra-session and inter-session similarities. The fiducial point timing variability is also higher, ranging from 12 to 31 ms for the SCG and GCG signals, in comparison to the intra-session variability of 1 to 8 ms (Table 3).
@@ 3.4.1. Motion artifact removal pipeline
We recorded a dataset containing six types of motion events: brow raising, chewing, nodding, talking, and walking. All recordings were collected from a single subject using in-ear microphones. This subject was a member of the study team who provided informed written consent prior to any data collection. For each acoustic event, we recorded 10-second segments, a duration that is in line with standard practice in large scale audio datasets such as AudioSet (Gemmeke et al., 2017), which uses this duration to capture sufficient acoustic context for a diverse range of events for use in machine learning models. This choice does not limit how frequently our system can output results because (1) a real-time implementation can use a sliding window with a small overlap between windows and (2) our waveform reconstruction model can run in real time and complete inference on a 10 second clip in 19 ms on a smartphone.
We also recorded a set of recordings when the user was static and when there was no external interference or music playback (Fig. 9a). Fig. 9b shows that the various motion types produced distinct signal distortions that were clearly visible and would obscure the smaller heart sounds.